# EVA-MILP: Towards Standardized Evaluation of MILP Instance Generation

### v2

@@ Abstract
Mixed-Integer Linear Programming (MILP) is fundamental to solving complex decision-making problems. The proliferation of MILP instance generation methods, driven by machine learning’s demand for diverse optimization datasets and the limitations of static benchmarks, has significantly outpaced standardized evaluation techniques. Consequently, assessing the fidelity and utility of synthetic MILP instances remains a critical, multifaceted challenge. This paper introduces a comprehensive benchmark framework designed for the systematic and objective evaluation of MILP instance generation methods. Our framework provides a unified and extensible methodology, assessing instance quality across crucial dimensions: mathematical validity, structural similarity, computational hardness, and utility in downstream machine learning tasks. A key innovation is its in-depth analysis of solver-internal features –particularly by comparing distributions of key solver outputs including root node gap, heuristic success rates, and cut plane usage – leveraging the solver’s dynamic solution behavior as an ‘expert assessment’ to reveal nuanced computational resemblances. By offering a structured approach with clearly defined solver-independent and solver-dependent metrics, our benchmark aims to facilitate robust comparisons among diverse generation techniques, spur the development of higher-quality instance generators, and ultimately enhance the reliability of research reliant on synthetic MILP data. The framework’s effectiveness in systematically comparing the fidelity of instance sets is demonstrated using contemporary generative models. The code is available in https://github.com/anonymous-neurips-submission-2025/EVA-MILP.
@@ 1 Introduction
Mixed-Integer Linear Programming (MILP) is a fundamental optimization framework extending standard Linear Programming (LP) by incorporating integer variables, enabling the modeling of discrete decisions and logical conditions often intractable with purely continuous models [1, 2]. Its versatility in capturing complex relationships like fixed costs, mutual exclusivity, and indivisible entities makes MILP essential for practical decision-making in fields such as supply chain optimization, scheduling, financial modeling, and network design [3, 4, 5, 6].
@@
Traditionally, MILP solver advancement relied on benchmark libraries like MIPLIB [7] and other real-world instance collections. However, these static benchmarks often lack the scale, diversity, or controllable characteristics vital for modern research. This limitation is particularly acute with the rise of machine learning (ML) in optimization, where learning-based approaches for tasks like branching strategy selection or algorithm configuration demand large, varied datasets that often exceed available collections [8]. Consequently, a significant shift towards proactive generation of synthetic MILP instances has occurred [9, 10, 11, 12, 13, 14, 15, 16]. This trend towards instance generation is motivated by multiple advantages: fulfilling the need for extensive datasets in ML applications; achieving greater instance diversity than found in existing libraries [7]; enabling control over computational difficulty for rigorous algorithm testing; facilitating the simulation of specific problem structures [13]; aiding solver testing and debugging [17]; and creating privacy-preserving surrogates for confidential real-world data.
@@
Methodologies for MILP instance generation have evolved in tandem. Early approaches featured parameterized generators for specific problem classes such as the traveling salesman problems [18, 19], set covering [20], or quadratic assignment problems [21]. Later techniques emphasized more generalized feature-based descriptions and sampling [9, 22]. Most recently, deep learning based frameworks like G2MILP [10], MILP-StuDio [13], and MILP-Evolve [23] have gained prominence. These often represent MILP instances as graphs (e.g., bipartite graphs) and use generative models like variational autoencoders [24] to learn and replicate specific structures from training data, marking a significant potential advancement but also introducing new evaluation challenges.
@@
Despite this progress in generation, a critical and multifaceted challenge remains: the fair, objective, and comprehensive evaluation of synthetic MILP instance “quality.” Assessing generator outputs effectively requires considering several dimensions: (I) Ensuring fundamental solvability (feasibility and boundedness) is crucial yet non-trivial [11, 10, 13]. (II) Generated instances should ideally mirror real-world structural patterns [7], but mere resemblance may not capture true computational behavior. (III) Instances must exhibit appropriate and controllable computational hardness, typically evaluated via solver performance [17], though achieving desired hardness levels consistently is a significant difficulty. Furthermore, current evaluation protocols [10, 9, 16] often lack standardization, hindering robust comparisons and impeding progress.
To address the aforementioned challenges and lack of standardization, this research introduces a novel benchmark framework for the systematic and comprehensive evaluation of MILP instance generation techniques. Our framework provides a unified, extensible methodology to assess instance quality along key dimensions – mathematical validity, structural similarity, computational hardness, solver interaction patterns, and utility for downstream ML tasks – aiming to establish an objective standard. This, in turn, will facilitate fairer comparisons, guide the development of higher-quality generators, and improve the reliability of research utilizing synthetic MILP data. Specifically, our contributions are: (I) We define and integrate key evaluation metrics into a unified framework, assessing instances based on both Solver-Independent Metrics (covering fundamental properties and structural resemblance) and Solver-Dependent Metrics (including computational characteristics and downstream task utility), thereby providing a more holistic assessment than existing evaluations. (II) Our framework uniquely incorporates detailed analysis of solver-internal features (e.g., root node gap, heuristic success counts, cut plane profiles analyzed via PCA and Wasserstein distance), treating the solver’s behavior as an expert assessment to reveal deeper computational similarities beyond surface structure. (III) The proposed benchmark is intentionally designed with extendability in mind; its modular structure facilitates the future integration of additional evaluation metrics, new datasets representing diverse MILP problem types, and novel generation techniques as the field progresses, ensuring its long-term relevance. (IV) It enables objective and fair comparisons between diverse MILP instance generators (including rule-based, statistical, and machine learning-based approaches) by offering clear, operational metrics and quantitative criteria, thereby providing guidance for the development of higher-quality generation methods for the community.
@@
Furthermore, this work highlights a limitation in current DL-based generation approaches. While representing MILP instances as bipartite graphs has enabled progress, particularly in achieving structural similarity, this representation may struggle to fully capture the intricate constraint relationships and inherent mathematical properties crucial for MILP feasibility and hardness. We posit that future advancements in MILP instance generation might benefit from shifting focus beyond direct graph structure manipulation (like node/constraint operations) towards methods that more directly model or preserve the underlying mathematical structure, or solution space characteristics. This proposed benchmark provides the necessary tools to evaluate such novel approaches rigorously.
@@ 2 Preliminaries
MILP and its Bipartite Graph Representation. We consider the standard Mixed-Integer Linear Programming (MILP) problem:
@@
Representing MILP instances as weighted bipartite graphs is an established practice in the relevant literature [25, 26, 27]. In this common representation, each MILP instance corresponds to a graph $\mathcal{G}$ defined by three sets: constraint vertices $\mathcal{C}$, variable vertices $\mathcal{V}$, and edges $\mathcal{E}$ connecting them. The constraint partition $\mathcal{C}=\{c_{1},\ldots,c_{m}\}$ includes one vertex $c_{i}$ for each of the $m$ constraints, where the vertex feature $c_{i}$ typically represents the bias term, i.e., $c_{i}=(b_{i})$ (the $i$-th element of $\mathbf{b}$ from the MILP formulation). The variable partition $\mathcal{V}=\{v_{1},\ldots,v_{n}\}$ contains one vertex $v_{j}$ for each of the $n$ variables, with the corresponding 9-dimensional feature vector $\mathbf{v}_{j}$ encoding information such as the objective coefficient $c_{j}$ (from the MILP vector $\mathbf{c}$), the variable type, and the bounds $l_{j},u_{j}$ (from the MILP vectors $\mathbf{l}$ and $\mathbf{u}$). Edges in $\mathcal{E}$ only exist between vertices from different partitions ($c_{i}\in\mathcal{C}$ and $v_{j}\in\mathcal{V}$). An edge $e_{i,j}$ connects $c_{i}$ and $v_{j}$ if the corresponding element $A_{i,j}$ in the constraint matrix $\mathbf{A}$ is non-zero; its associated feature $e_{i,j}$ is described by this coefficient, i.e., $e_{i,j}=(A_{i,j})$. If $A_{i,j}$ is zero, the edge $e_{i,j}$ is considered absent.
@@ 3 Methodology
This section outlines the methodology employed to evaluate the quality and characteristics of generated MILP instances. We compare instances generated using reproductions of three open-source methods – G2MILP [10], ACM-MILP [12], and DIG-MILP [11] – against a baseline set of ‘original’ instances. This baseline set, comprising Set Cover, Combinatorial Auction (CA), Capacitated Facility Location (CFL), and Independent Set (IS) problems, was synthesized using the Ecole library [28] (the specific parameter settings used for this synthesis are detailed in the Appendix A), while IP and LB come from two challenging real-world problem families used in ML4CO 2021 competition [29]. To capture different facets of similarity and fidelity between the generated and original sets, we utilize a suite of metrics. These metrics are organized into two primary categories based on their reliance on the underlying optimization solver: (I) Solver-Independent Metrics, which assess inherent properties of the MILP instances themselves (e.g., feasibility, structure) without regard to any specific solution process; and (II) Solver-Dependent Metrics, which evaluate aspects intrinsically linked to the behavior and performance of a particular solver (primarily Gurobi) when applied to the instances. This dual approach allows for a robust assessment, considering both fundamental instance properties and their practical implications for optimization algorithms (See the whole framework in Figure 1).
Figure 1: Architecture of EVA-MILP
3.1 Dataset
This study evaluates MILP instances from two sources: original instances (baselines and validation) and those from three generative models (G2MILP, ACM-MILP, DIG-MILP). Baselines include four synthetic datasets (SC, CA, CFL, IS) from Ecole [28] (details in Appendix A) and public benchmarks from the ML4CO Competition 2024 [29] for pipeline validation. Due to the inherent specialization of generative models to particular problem types, our comparative analysis, which forms the core of our evaluation, specifically examines SC instances from G2MILP, CA instances from ACM-MILP and DIG-MILP, and IS instances from ACM-MILP and G2MILP to demonstrate our metrics. The methodology’s applicability to other datasets and models has been confirmed.
@@ 3.2.1 Feasibility Ratio
This metric measures the percentage of instances within a given dataset that are both feasible and bounded.
Results The feasibility and boundedness of generated instance sets for IS, CA, and SC problem domains were evaluated. Most generative models successfully maintained 100% feasibility and boundedness, mirroring the original datasets. A minor exception was G2MILP (MIS, $\eta=0.10$), which had a 93.40% feasibility ratio. This suggests that while generally robust, some model configurations, particularly with specific problem structures and higher mask ratios, might occasionally produce invalid instances. Results could be found in Table 10, 11, and 12 in the Appendix.
@@ 3.2.2 Structural Similarity
To quantitatively assess the structural similarity between two sets of MILP instances, following [10], we implemented a feature-based comparison method. Each MILP instance is first converted into a graph representation. From this representation, a vector of 11 predefined structural features is computed for each instance. For each of the 11 features, we compute the Jensen-Shannon Divergence (JSD) between the distributions observed in the two sets. The JSD value for each feature $i$ is transformed into a similarity score $S_{i}$ using the formula $S_{i}=1-\frac{JSD_{i}}{\log(2)}$, where a score of 1 indicates identical distributions and 0 indicates maximal divergence. The overall structural similarity score between the two sets is then calculated as the arithmetic mean of the individual feature similarity scores ($S_{i}$).
Table 1: Structural Similarity Across Problem Types, Models, and Mask Ratios ($\eta$)
Problem Model Overall Similarity at Masked Ratio ($\eta$)
@@
IS G2MILP 0.486 0.473 0.466
SC G2MILP 0.990 0.950 0.921
Results The JSD-based metric (Table 1) shows that models vary in their ability to replicate structural characteristics across problem domains. For CA instances, ACM-MILP and DIG-MILP achieved high structural similarity to reference data (scores 0.860-0.892). In contrast, for IS problems, ACM-MILP showed moderate similarity (0.699-0.734), while G2MILP’s scores were lower (0.466-0.486). G2MILP, however, excelled with Set Cover problems, yielding very high similarity (0.921-0.990), particularly at lower mask ratios.
@@ 3.3.1 Branching nodes
Computational hardness was assessed using the number of branch-and-bound nodes [30] explored by the Gurobi solver during the optimization process. The formula used is:
Relative $\text{Relative Error}=\left|\frac{\sum Nodes_{\text{generated}}-\sum Nodes_{\text{training}}}{\sum Nodes_{\text{training}}}\right|\times 100\%$
Results The Gurobi branching node counts (Appendix. F Table 13) reveal that computational hardness is highly sensitive to both the generative model and problem type. For IS instances, G2MILP drastically increased complexity (e.g., RE > 50,000% for $\eta=0.10$), leading to frequent time-outs, while ACM-MILP instances remained close to the baseline’s simplicity. A similar divergence occurred with CA instances: ACM-MILP produced substantially harder problems than the baseline (e.g., RE > 26,000% for $\eta=0.20$), whereas DIG-MILP only modestly increased difficulty. For SC problems, G2MILP’s generated instances were moderately harder but remained tractable.

### v3

# Are Your Generated Instances Truly Useful? GenBench-MILP: A Benchmark Suite for MILP Instance Generation

@@ Abstract
The proliferation of machine learning-based methods for Mixed-Integer Linear Programming (MILP) instance generation has surged, driven by the need for diverse training datasets. However, a critical question remains: Are these generated instances truly useful and realistic? Current evaluation protocols often rely on superficial structural metrics or simple solvability checks, which frequently fail to capture the true computational complexity of real-world problems. To bridge this gap, we introduce GenBench-MILP, a comprehensive benchmark suite designed for the standardized and objective evaluation of MILP generators. Our framework assesses instance quality across four key dimensions: mathematical validity, structural similarity, computational hardness, and utility in downstream tasks. A distinctive innovation of GenBench-MILP is the analysis of solver-internal features—including root node gaps, heuristic success rates, and cut plane usage. By treating the solver’s dynamic behavior as an expert assessment, we reveal nuanced computational discrepancies that static graph features miss. Our experiments on instance generative models demonstrate that instances with high structural similarity scores can still exhibit drastically divergent solver interactions and difficulty levels. By providing this multifaceted evaluation toolkit, GenBench-MILP aims to facilitate rigorous comparisons and guide the development of high-fidelity instance generators.
Machine Learning, ICML
{NoHyper}
@@ 1 Introduction
Mixed-Integer Linear Programming (MILP) is a fundamental optimization framework extending standard Linear Programming (LP) by incorporating integer variables, enabling the modeling of discrete decisions and logical conditions often intractable with purely continuous models (Achterberg and Wunderling, 2013; Wolsey, 2020). Its versatility in capturing complex relationships like fixed costs, mutual exclusivity, and indivisible entities makes MILP essential for practical decision-making in fields such as supply chain optimization, scheduling, financial modeling, and network design (Hugos, 2018; Branke et al., 2015; Mansini et al., 2015; Al-Falahy and Alani, 2017).
@@
Traditionally, MILP solver advancement relied on benchmark libraries like MIPLIB (Gleixner et al., 2021) and other real-world instance collections. However, these static benchmarks often lack the scalability, diversity, or controllable characteristics vital for modern research. These limitations are particularly acute with the rise of machine learning (ML) in optimization, where learning-based approaches for tasks like branching strategy selection or algorithm configuration demand large, varied datasets that often exceed available collections (Bengio et al., 2021). Consequently, a significant shift towards proactive generation of MILP instances has occurred (Bowly and others, 2020; Geng and others, 2023; Wang and others, 2023; Guo and others, 2024; Liu and others, 2024; Yang et al., 2024b; Zeng and others, 2024; Zhang and others, 2024). This trend towards instance generation is motivated by multiple highlighted advantages: fulfilling the need for extensive datasets in ML applications; achieving greater instance diversity than found in existing libraries (Gleixner et al., 2021); enabling control over computational difficulty for rigorous algorithm testing; facilitating the simulation of specific problem structures (Liu and others, 2024); aiding solver testing and debugging (Gurobi Optimization LLC, 2021); and creating privacy-preserving surrogates for confidential real-world data.
Figure 1: Architecture of GenBench-MILP
@@
Methodologies for MILP instance generation have evolved in tandem. Early approaches featured parameterized generators for specific problem classes such as the traveling salesman problems (Pilcher and Rardin, 1992; Vander Wiel and Sahinidis, 1995), set covering (Balas and Ho, 1980), or quadratic assignment problems (Culberson, 2002). Later techniques emphasized more generalized feature-based descriptions and sampling (Bowly and others, 2020; Smith-Miles and Bowly, 2015). Most recently, deep learning based frameworks like G2MILP (Geng and others, 2023), MILP-StuDio (Liu and others, 2024), and MILP-Evolve (Li and others, 2024) have gained prominence. These often represent MILP instances as graphs (e.g., bipartite graphs) and use generative models like variational autoencoders (Kipf and Welling, 2016) to learn and replicate specific structures from training data, marking a significant potential advancement but also introducing new evaluation challenges. More information about related work could be found in Appendix B
@@
Despite this progress in generation, a critical and multifaceted challenge remains: the fair, objective, and comprehensive evaluation of synthetic MILP instance “quality”. Assessing generator outputs effectively requires considering several dimensions: (I) Ensuring fundamental solvability (feasibility and boundedness) is crucial yet non-trivial (Wang and others, 2023; Geng and others, 2023; Liu and others, 2024). (II) Generated instances should ideally mirror real-world structural patterns (Gleixner et al., 2021), but mere resemblance may not capture true computational behavior. (III) Instances must exhibit appropriate and controllable computational hardness, typically evaluated via solver performance (Gurobi Optimization LLC, 2021), though achieving desired hardness levels consistently is a significant difficulty. Furthermore, current evaluation protocols (Geng and others, 2023; Bowly and others, 2020; Zhang and others, 2024) often lack standardization, hindering robust comparisons and impeding progress.
@@
To address the lack of a standardized and comprehensive methodology for evaluating MILP instance generation techniques, this research introduces a novel benchmark framework. Our unified and extensible framework assesses instance quality across key dimensions – mathematical validity, structural similarity, computational hardness, solver interaction patterns, and utility for downstream ML tasks. By establishing an objective standard, we aim to facilitate fairer comparisons, guide the development of higher-quality generators, and improve the reliability of research utilizing synthetic MILP data. The practical viability and robustness of our framework is comfirmed by our tests on the super hard dataset with a strict 120-second time limit, proving our solver-internal features are highly stable and confirming the approach is both low-cost and efficient. Our experiments on open-source solvers SCIP and HiGHS further confirm that while the framework is general, state-of-the-art solvers like Gurobi yield more precise internal features. These findings not only validate our framework but also highlight a limitation in current DL-based generation approaches. While representing MILP instances as bipartite graphs has enabled progress in achieving structural similarity, this representation may struggle to fully capture the intricate constraint relationships and inherent mathematical properties crucial for MILP feasibility and hardness. We posit that future advancements might benefit from shifting focus from direct graph structure manipulation towards methods that more directly model the underlying mathematical structure or solution space characteristics. The proposed benchmark provides the necessary tools to rigorously evaluate such novel approaches.
Our specific contributions are four-fold: (I) We integrate both Solver-Independent Metrics (covering fundamental properties and structural resemblance) and Solver-Dependent Metrics (including computational characteristics and downstream task utility) for a holistic assessment. (II) We incorporate detailed analysis of solver-internal features (e.g., root node gap, heuristic success counts, cut plane profiles), using the solver’s behavior as an expert assessment to reveal deeper computational similarities. (III) The framework is designed for extendability, featuring a modular structure to easily accommodate new metrics, datasets, and generation techniques over time. (IV) We provide the community with clear, operational metrics and quantitative criteria to objectively benchmark diverse generator types (e.g., rule-based, statistical, and machine learning-based), thereby steering future development.
@@ 2 Preliminaries
MILP and its Bipartite Graph Representation. We consider the standard MILP problem:
@@
Representing MILP instances as weighted bipartite graphs is an established practice in the relevant literature (Zhang et al., 2023; Gasse et al., 2019; Nair et al., 2021). In this common representation, each MILP instance corresponds to a graph $\mathcal{G}$ defined by three sets: constraint vertices $\mathcal{C}$, variable vertices $\mathcal{V}$, and edges $\mathcal{E}$ connecting them. The constraint partition $\mathcal{C}=\{c_{1},\ldots,c_{m}\}$ includes one vertex $c_{i}$ for each of the $m$ constraints, where the vertex feature $c_{i}$ typically represents the bias term, i.e., $c_{i}=(b_{i})$ (the $i$-th element of $\mathbf{b}$ from the MILP formulation). The variable partition $\mathcal{V}=\{v_{1},\ldots,v_{n}\}$ contains one vertex $v_{j}$ for each of the $n$ variables, with the corresponding 9-dimensional feature vector $\mathbf{v}_{j}$ encoding information such as the objective coefficient $c_{j}$ (from the MILP vector $\mathbf{c}$), the variable type, and the bounds $l_{j},u_{j}$ (from the MILP vectors $\mathbf{l}$ and $\mathbf{u}$). Edges in $\mathcal{E}$ only exist between vertices from different partitions ($c_{i}\in\mathcal{C}$ and $v_{j}\in\mathcal{V}$). An edge $e_{i,j}$ connects $c_{i}$ and $v_{j}$ if the corresponding element $A_{i,j}$ in the constraint matrix $\mathbf{A}$ is non-zero; its associated feature $e_{i,j}$ is described by this coefficient, i.e., $e_{i,j}=(A_{i,j})$. If $A_{i,j}$ is zero, the edge $e_{i,j}$ is considered absent.
@@
Feasibility  Consider the MILP problem defined by constraints $\mathbf{A}\mathbf{x}\leq\mathbf{b}$ and $x_{i}\in\mathbb{Z},\forall i\in\mathcal{I}$. The feasible region is the set $\mathcal{F}=\{\mathbf{x}\in\mathbb{R}^{n}\mid\mathbf{A}\mathbf{x}\leq\mathbf{b},x_{i}\in\mathbb{Z}\;\forall i\in\mathcal{I}\}$. The MILP problem is feasible if its feasible region $\mathcal{F}$ is non-empty, i.e., $\mathcal{F}\neq\emptyset$. A specific point $\hat{\mathbf{x}}\in\mathbb{R}^{n}$ is considered feasible if it satisfies all constraints, i.e., $\hat{\mathbf{x}}\in\mathcal{F}$.
@@ 3 Methodology
This section outlines the methodology employed to evaluate the quality and characteristics of generated MILP instances. We compare instances generated using reproductions of three open-source methods – G2MILP (Geng and others, 2023), ACM-MILP (Guo and others, 2024), and DIG-MILP (Wang and others, 2023) – against a baseline set of ‘original’ instances. This baseline set, comprising Set Cover, Combinatorial Auction (CA), Capacitated Facility Location (CFL), and Independent Set (IS) problems, was synthesized using the Ecole library (Prouvost et al., 2020) (the specific parameter settings used for this synthesis are detailed in the Appendix A.1), while IP and LB come from two challenging real-world problem families used in ML4CO 2021 competition (Gasse et al., 2022). To capture different facets of similarity and fidelity between the generated and original sets, we utilize a suite of metrics. These metrics are organized into two primary categories based on their reliance on the underlying optimization solver: (I) Solver-Independent Metrics, which assess inherent properties of the MILP instances themselves (e.g., feasibility, structure) without regard to any specific solution process; and (II) Solver-Dependent Metrics, which evaluate aspects intrinsically linked to the behavior and performance of a particular solver (primarily Gurobi) when applied to the instances. This dual approach allows for a robust assessment, considering both fundamental instance properties and their practical implications for optimization algorithms (See the whole framework in Figure 1).
3.1 Datasets
This study evaluates MILP instances from two sources: original instances (baselines and validation) and those from three generative models (G2MILP, ACM-MILP, DIG-MILP). Baselines include four synthetic datasets (SC, CA, CFL, IS) from Ecole (Prouvost et al., 2020) (details in Appendix A.1) and public benchmarks from the ML4CO Competition 2024 (Gasse et al., 2022) for pipeline validation. Due to the inherent specialization of current generative models to particular problem types, our comparative analysis specifically examines SC instances from G2MILP, CA instances from ACM-MILP and DIG-MILP, and IS instances from ACM-MILP and G2MILP. We emphasize that these specific combinations serve primarily as representative examples to demonstrate the utility of our metrics; the GenBench-MILP framework itself is dataset-agnostic, and its broader applicability to other distinct datasets and models has been confirmed (see Section 4.1).
@@ 3.2.1 Feasibility Ratio
This metric measures the proportion of instances that are both feasible and bounded. Most methods maintain near-perfect feasibility across datasets, with only minor exceptions (e.g., G2MILP on MIS at $\eta=0.10$). Detailed results are given in Table 12 in the Appendix.
@@ 3.2.2 Structural Similarity
Table 1: Structural Similarity. The similarity score is derived from Jensen-Shannon Divergence (higher is better). The Masked Ratio ($\eta$) denotes the proportion of the original instance’s structure (e.g., constraint nodes in G2MILP or constraints in ACM-MILP) that is masked, corrupted, or selected for reconstruction during the generation process, serving as a control parameter for the modification magnitude.
Problem Model Similarity at Masked Ratio ($\eta$)
@@
[…]
To quantitatively assess the structural similarity between two sets of MILP instances, following (Geng and others, 2023), we implemented a feature-based comparison method. Each instance is first converted into a graph representation, from which a vector of 11 predefined structural features is extracted. For each feature, we compute the Jensen–Shannon Divergence (JSD) between the distributions of the two sets and transform it into a similarity score $S_{i}=1-\frac{JSD_{i}}{\log(2)}$, where a score of 1 indicates identical distributions and 0 indicates maximal divergence. The overall structural similarity between the two sets is the arithmetic mean of these feature scores. Overall, the JSD-based metric (Table 1) indicates that ACM-MILP and DIG-MILP closely match the structural characteristics of CA data (around 0.86–0.89), while ACM-MILP achieves moderate similarity on IS (around 0.70) and G2MILP is lower (around 0.47). For Set Cover, G2MILP performs best, reaching very high similarity (around 0.92–0.99), especially at lower mask ratios.
@@ 3.3.1 Branching nodes
[…]
Table 2: Branching Node Statistics and Relative Error across Problem Types. Statistics are calculated over 1000 instances per source/ratio. Original refers to the original training sets (Total Nodes: IS=14,255; CA=18,488; SC=46,408). ’Time Limit Hits’ indicates premature termination. Relative Error (RE) compares generated total nodes to baseline total nodes. $\eta$ denotes the mask ratio.
Problem Source Mask Ratio ($\eta$) Mean Nodes Median Nodes Std Dev Max Nodes Time Limit Hits Relative Error (%)
IS Original N/A 14.3 1.0 49.1 1,056 0
G2MILP 0.01 26.2 1.0 60.1 724 0 84.1
G2MILP 0.05 765.7 175.0 1,941.6 16,314 5 5,271.6
[…]
ACM-MILP 0.05 8.8 1.0 36.8 572 0 38.3
[…]
CA Original N/A 18.5 1.0 48.1 514 0
[…]
ACM-MILP 0.20 4,873.6 962.5 8,021.7 43,036 75 26,259.8
DIG-MILP 0.01 34.1 1.0 76.2 799 0 84.4
DIG-MILP 0.05 34.3 1.0 76.4 812 0 85.6
DIG-MILP 0.10 37.8 1.0 83.5 902 0 104.3
SC Original N/A 46.4 1.0 298.6 8,135 0
G2MILP 0.01 85.4 1.0 536.5 14,592 0 84.0
G2MILP 0.05 79.6 1.0 430.9 10,643 0 71.6
[…]